I am trying to come up with an algebraic equation to determine the cost of two refrigerators with good energy ratings. I found one with that will cost .1143 cents per kWh and it boasts that it uses 486 kWh per year. I want to compare it with another one I found that will cost .1143 cents per kWh. This one uses 526 kWh per year. If the purchase price of the first refrigerator is $950 and the second is $875, which one will be the better buy after 10 years?
I'm sorry for the long question, I hope someone out there can give me an algebraic equation for putting this all together. I'd like to use it again for other appliances in the future.''.

Answer by mananth(16946)

First refrigerator cost + power cost for ten years
General Equation:
Let the number of years run be x
(Cx1)=950+0.1143*486*x
(Cx)=950 + 55.55x
Second refrigerator
(Cx2)=875+0.1143*526*x
(Cx2)=875+60.12x
...
Plug the value of x=10 for both equations to find which is a better deal.
(Cx1)= $1505.5
(Cx2)= $ 1476.2
..
You decide
..
m.ananth@hotmail.ca